What pressure (in atm) would be exerted by a mixture of 1.4 g of Chlorine gas (Mm= 70 g/mol) and 4.8 g of nitrogen gas (Mm= 28 g/mole) in a 200 mL container at 57oC?


1
Expert's answer
2021-12-16T11:15:03-0500

Moles of chlorine "\frac{1.4}{70}=0.02moles"

Moles of nitrogen "\frac{4.8}{28}=0.17moles"


PV= nRT


"P\u00d70.2=[0.02+0.17]\u00d70.0821\u00d7[57+273]"


"P=\frac{0.19×0.0821×330}{0.2}=25.73atm"